# Bregman Stochastic Proximal Point Algorithm with Variance Reduction

### v1

@@ Abstract
Stochastic algorithms, especially stochastic gradient descent (SGD), have proven to be the go-to methods in data science and machine learning. In recent years, the stochastic proximal point algorithm (SPPA) emerged, and it was shown to be more robust than SGD with respect to stepsize settings. However, SPPA still suffers from a decreased convergence rate due to the need for vanishing stepsizes, which is resolved by using variance reduction methods. In the deterministic setting, there are many problems that can be solved more efficiently when viewing them in a non-Euclidean geometry using Bregman distances. This paper combines these two worlds and proposes variance reduction techniques for the Bregman stochastic proximal point algorithm (BSPPA). As special cases, we obtain SAGA- and SVRG-like variance reduction techniques for BSPPA. Our theoretical and numerical results demonstrate improved stability and convergence rates compared to the vanilla BSPPA with constant and vanishing stepsizes, respectively. Our analysis, also, allow to recover the same variance reduction techniques for Bregman SGD in a unified way.
@@ 1 Introduction
Contributions. In this paper we propose a Bregman Stochastic Proximal Point Algorithm (BSPPA) with generic variance reduction. We provide improved convergence rates as compared with vanilla BSPPA without variance reduction, in particular, sublinear and linear rates for convex or relatively strongly convex functions, respectively. Then, from the generic results, convergence rates for BSAPA, BSVRP, and L-BSVRP, which are proximal and Bregman distance versions of SAGA [9], SVRG [15], and L-SVRG [20] respectively. These variance-reduced algorithms are all new in the literature. We can also recover rates for vanilla BSPPA for potentially non-relatively smooth functions. Of course, the rates of the variance-reduced versions (BSAPA, BSVRP, L-BSVRP) are given without the need for vanishing stepsizes, and are faster than that of BSPPA, which are the intended objectives of variance reduction. As a by-product of our study, we provide a unified variance reduction analysis for Bregman SGD. Our analysis can thus recover not only the Bregman version of SAGA and L-SVRG (present in Dragomir et al. [11]), but also Bregman SVRG and the Bregman version of the unified study in Gorbunov et al. [13], dealing with several variants of (variance-reduced) Euclidean SGD.
@@ 2.1 Notation
Let $\mathcal{C}\subset\mathbb{R}^{d}$ be convex with non empty interior. We denote interior of $\mathcal{C}$ by $\operatorname*{int}\mathcal{C}$ and its closure by $\bar{\mathcal{C}}$. For every integer $\ell\geq 1$, we define $\ell]:=\{1,\dots,\ell\}$. Bold default font is used for random variables taking values in $\mathbb{R}^{d}$, while bold sans serif font is used for their realizations or deterministic variables in $\mathbb{R}^{d}$. The probability space underlying random variables is denoted by $(\Omega,\mathfrak{A},\mathsf{P})$. For every random variable x{x}, $\mathsf{E}[{x}]$ denotes its expectation, while if $\mathfrak{F}\subset\mathfrak{A}$ is a sub $\sigma$-algebra we denote by $\mathsf{E}[{x}\,|\,\mathfrak{F}]$ the conditional expectation of x{x} given $\mathfrak{F}$. Also, $\sigma({y})$ represents the $\sigma$-algebra generated by the random variable y{y}. For a function ${\varphi}\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$, define $\operatorname*{dom}{\varphi}\coloneqq\left\{{\mathsf{x}}\in\mathbb{R}^{d}:{\varphi}({\mathsf{x}})<+\infty\right\}$. $\varphi$} is proper if $\operatorname*{dom}\varphi\neq\varnothing$. The set of minimizers of $\varphi$} is $\operatorname*{arg\,min}{\varphi}=\{{\mathsf{x}}\in\mathbb{R}^{d}:{\varphi}({\mathsf{x}})=\inf{\varphi}\}$. If $\inf{\varphi}$ is finite, it is represented by $\varphi$$_{*}$. We denote the subdifferential of $\varphi$ at $\mathsf{x}$} as $\partial\varphi({\mathsf{x}})\coloneqq\left\{{{\mathsf{g}}}\in\mathbb{R}^{d}:{\varphi}({\mathsf{y}})-{\varphi}({\mathsf{x}})\geq\langle{{\mathsf{g}}},{\mathsf{y}}-{\mathsf{x}}\rangle\;(\forall\,{\mathsf{y}}\in\mathbb{R}^{d})\right\}$. When $\varphi$} is differentiable $\nabla{\varphi}$ denotes the gradient of $\varphi$}. The (convex) conjugate of $\varphi$} is the function ${\varphi}^{*}\colon\mathbb{R}^{d}\to[-\infty,+\infty]$ defined by ${\mathsf{y}}\mapsto\sup_{{\mathsf{x}}\in\mathbb{R}^{d}}\langle{\mathsf{y}},{\mathsf{x}}\rangle-{\varphi}({\mathsf{x}}).$ In this work, $\ell^{1}$ represents the space of sequences with summable norms and $\ell^{2}$ with summable squared norms.
@@ 2.2 Bregman distance
The Bregman distance $D_{h}({\mathsf{x}},{\mathsf{y}})$, between ${\mathsf{x}}\in\mathcal{C},{\mathsf{y}}\in\operatorname{int}\mathcal{C}$, is defined as $D_{h}({\mathsf{x}},{\mathsf{y}})\coloneqq h({\mathsf{x}})-h({\mathsf{y}})-\langle\nabla h({\mathsf{y}}),{\mathsf{x}}-{\mathsf{y}}\rangle$, where $h\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$, often called kernel, is a strictly convex and twice continuously differentiable function on $\operatorname{int}\mathcal{C}$, with $\operatorname*{dom}h=\mathcal{C}$. In what follows, $h$ will always refer to a function with those properties, unless stated otherwise. A classical example is the usual squared Euclidean distance $D_{h}({\mathsf{x}},{\mathsf{y}})=\frac{1}{2}\|{\mathsf{x}}-{\mathsf{y}}\|_{2}^{2}$ when $h=\frac{1}{2}\|\cdot\|_{2}^{2}$. The Bregman distance often captures the underlying geometry of an optimization problem more effectively than the Euclidean distance. It can be better suited to the constraints set or to the objective function properties. This makes it a powerful tool in optimization. For instance, the Kullback–Leibler divergence on $\mathbb{R}^{d}$ does not have a global Lipschitz continuous gradient, however it is relatively smooth w.r.t. the kernel $h({\mathsf{x}})=-\sum_{i}\log\mathsf{x}_{i}$; see Bauschke et al. [3]. Unfortunately, the Bregman distance is not necessarily symmetric, homogeneous nor translation invariant. The latter two shortcomings are the reasons why the following assumption is required.
Assumptions 2.1.
@@
$D_{h^{*}}({\mathsf{x}}+\lambda{\mathsf{z}},{\mathsf{x}})\leq G({\mathsf{x}},{\mathsf{y}},{\mathsf{z}})\lambda^{2}D_{h^{*}}({\mathsf{y}}+{\mathsf{z}},{\mathsf{y}}),$ (2.1)
According to Dragomir et al. [11], this assumption seems unavoidable when using past iterates in an algorithm with Bregman distance. This is also the case of accelerated methods where similar assumption is made; see Hanzely et al. [14].
@@
    From Bauschke et al. [2, Theorem 3.7], we know that $D_{h}({\mathsf{x}},{\mathsf{y}})=D_{h^{*}}(\nabla h({\mathsf{y}}),\nabla h({\mathsf{x}}))$ for all $({\mathsf{x}},{\mathsf{y}})\in(\operatorname{int}\,\mathcal{C})^{2}$. Hence, $\gamma_{h}=\gamma_{h^{*}}$.
@@ 3 Problem setting
First, we (re)introduce the finite-sum problem and collect all assumptions that we need for this paper, while not all of them will be required to hold at the same time. Section 4 presents the proposed algorithm for solving such problems:
@@
    ${\mathsf{x}}_{*}\in\operatorname*{int}\mathcal{C}$ such that $\nabla F({\mathsf{x}}_{*})=0$.
Assumption (A.i) is always supposed true in this paper. Assumptions (A.ii) and (A.iv) extend the Euclidean notions of strong convexity and $L$-smoothness (i.e., $L$-Lipschitz continuity of the gradient), respectively, to the Bregman setting. Assumption (A.ii), with $\beta>0$, is only required when establishing a linear convergence rate (to a neighborhood of the optimum) for the vanilla BSPPA algorithm. Similarly, (A.iv), which corresponds to the relative strong convexity of $F$, will be used to derive linear convergence rates to the exact minimum for the variance-reduced variants of BSPPA; in that case, only convexity of each $f_{i}$ will be assumed. Finally, (A.iv) and (A.ii) with $\beta>0$ are not required to obtain the sublinear convergence rates. Regarding Assumption (A.v), our main focus is not on the constrained cases where it may fail. Instead, we aim to address, possibly unconstrained, problems whose objective functions lack Lipschitz-continuous gradients—rendering standard Euclidean algorithms unsuitable—but are relatively smooth with respect to a Bregman kernel. Assumption (A.v) will be stated whenever needed.
@@ 4 BSPPA with generic variance reduction
We propose to solve Problem (3.1) using the following Bregman stochastic proximal point algorithm (BSPPA) with a generic variance reduction satisfying the abstract conditions in Assumption 4.5 further below.
@@
We assume in all this work that this minimization is well-posed, meaning there exists a unique solution in $\operatorname*{int}\mathcal{C}$. Compared to BSPPA, we have an additional linear perturbation that contains ${e}_{k}$. It is the generic variance reduction term. As we shall see in Section 5, depending on the specific algorithms, ${e}_{k}$ may be defined in various ways. When ${e}_{k}$ is set to 0, BSPPA is recovered. Using the optimality condition, the update at iteration $k$ can be rewritten as:
@@
where ${g}_{k+1}\in\partial f_{i_{k}}({x}_{k+1})$ is such that (4.1) holds. Since ${w}_{k}$ depends on ${g}_{k+1}$, Equation (4.1) shows that Algorithm 4.1 is an implicit algorithm. This is in contrast to an explicit update
@@
and is assumed well-posed, i.e., a unique solution in $\operatorname*{int}\mathcal{C}$ exists for any $k\in\mathbb{N}$.
@@
The virtual explicit sequences $({v}_{k})_{k\in\mathbb{N}}$ and $({z}_{k})_{k\in\mathbb{N}}$ introduced in (4.2) and (4.3) are crucial to our analysis. They will be used to prove the main proposition even though they do not appear in Algorithm 4.1.
@@
The integration of these explicit iterates is the reason why our analysis extends seamlessly to the Bregman SGD case, providing a unified variance reduction study for Bregman SGD. All the variance reduction results will stand true for SGD with Bregman distance; see Remark B.2 and Proposition B.3.
The next assumptions that we consider concern the noise and variance of Algorithm 4.1. The virtual explicit sequence $({z}_{k})_{k\in\mathbb{N}}$ appears in those assumptions for the sake of analysis. This allows, at the same time, the application of the analysis and the results to the Bregman SGD case.
@@
These assumptions on the noise and variance of the generic Algorithm 4.1 are “assumed” for the general analysis. But when algorithms are specified in Section 5 by defining ${e}_{k}$, they will be proved by corresponding lemmas. They read as follows.
@@
    $\mathsf{E}[{e}_{k}\,|\,\mathfrak{F}_{k}]=0$ a.s. and
    $\mathsf{E}[{g}_{k}\,|\,\mathfrak{F}_{k}]\in\partial F({x}_{k})$  a.s.,
@@
    $\mathsf{E}[D_{h}({x}_{k},{z}_{k+1})\,|\,\mathfrak{F}_{k}]\leq\alpha_{k}^{2}A_{k}\left(F({x}_{k})-F({\mathsf{x}}_{*})\right)+\alpha_{k}^{2}B_{k}\sigma_{k}^{2}+\alpha_{k}^{2}N_{k}$  a.s.,
@@
In the smooth case, Assumption (B.i) ensures that $\mathsf{E}[{v}_{k}\,|\,\mathfrak{F}_{k}]=\mathsf{E}[\nabla f_{i_{k}}({x}_{k})\,|\,\mathfrak{F}_{k}]=\nabla F({x}_{k})$, so that the direction ${v}_{k}$ is an unbiased estimator of the full gradient of $F$ at ${x}_{k}$, which is a standard assumption in the related literature. Assumption (B.ii) on $\mathsf{E}\left[D_{h}({x}_{k},{z}_{k+1})\,|\,\mathfrak{F}_{k}\right]$ is the equivalent of what is called, in the literature [17, 10] and in the Euclidean setting, the expected smoothness or $ABC$-assumption on $\mathsf{E}\left[\lVert\nabla f_{i_{k}}({x}_{k})\rVert^{2}\,|\,\mathfrak{F}_{k}\right]$ with $\sigma_{k}=\|\nabla F({x}_{k})\|$ and $N_{k}$ constant. Assumption (B.iii) provides some control on the variance from iteration to iteration. Indeed, as we will see later that the sequence $(\sigma_{k})_{k\in\mathbb{N}}$ encodes the variance of the Algorithm 4.1. Depending on $\rho$, (B.iii) makes sure that the variance does not blow up and possibly reduces along the iterations whenever the algorithm converges.

### v2

@@ Abstract
Stochastic algorithms, in particular stochastic gradient descent (SGD), have become the methods of choice in data science and machine learning. More recently, the stochastic proximal point algorithm (SPPA) has been introduced and shown to be more robust than SGD with respect to stepsize selection. However, SPPA still suffers from reduced convergence rates due to the need for vanishing stepsizes, a limitation that can be alleviated by variance reduction techniques. In the deterministic setting, many optimization problems can be solved more efficiently by exploiting non-Euclidean geometries through the use of Bregman distances. In this work, we bridge these two lines of research and propose variance reduction techniques for the Bregman stochastic proximal point algorithm (BSPPA). As special cases, our framework yields SAGA- and SVRG-type variance reduction methods for BSPPA. Our theoretical analysis and numerical experiments demonstrate improved stability and faster convergence rates compared to the vanilla BSPPA, both with constant and vanishing stepsizes, respectively. Moreover, our analysis also encompasses the SGD setting, allowing us to recover the same variance reduction techniques for Bregman SGD within a unified framework.
@@ 1 Introduction
Contributions. In this paper we propose a Bregman Stochastic Proximal Point Algorithm (BSPPA) with generic variance reduction. We provide improved convergence rates as compared with vanilla BSPPA without variance reduction, in particular, sublinear and linear rates for convex or relatively strongly convex functions, respectively. Then, from the generic results, convergence rates for BSAPA, BSVRP, and L-BSVRP, which are proximal and Bregman distance versions of SAGA [8], SVRG [15], and L-SVRG [20] respectively. These variance-reduced algorithms are all new in the literature. We can also recover rates for vanilla BSPPA for potentially nonsmooth functions. Of course, the rates of the variance-reduced versions (BSAPA, BSVRP, L-BSVRP) are given without the need for vanishing stepsizes, and are faster than that of BSPPA, which are the intended objectives of variance reduction. As a by-product of our study, we provide a unified variance reduction analysis for Bregman SGD. Our analysis can thus recover not only the Bregman version of SAGA and L-SVRG (present in Dragomir et al. [11]), but also Bregman SVRG and the Bregman version of the unified study in Gorbunov et al. [13], dealing with several variants of (variance-reduced) Euclidean SGD.
@@ 2.1 Notation
Let $\mathcal{C}\subset\mathbb{R}^{d}$ be convex with nonempty interior. We denote interior of $\mathcal{C}$ by $\operatorname*{int}\mathcal{C}$, its closure by $\bar{\mathcal{C}}$ and its boundary by $\text{bd}\,\mathcal{C}$. For every integer $\ell\geq 1$, we define $\ell]:=\{1,\dots,\ell\}$. Bold default font is used for random variables taking values in $\mathbb{R}^{d}$, while bold sans serif font is used for their realizations or deterministic variables in $\mathbb{R}^{d}$. The probability space underlying random variables is denoted by $(\Omega,\mathfrak{A},\mathsf{P})$. For every random variable x{x}, $\mathsf{E}[{x}]$ denotes its expectation, while if $\mathfrak{F}\subset\mathfrak{A}$ is a sub $\sigma$-algebra we denote by $\mathsf{E}[{x}\,|\,\mathfrak{F}]$ the conditional expectation of x{x} given $\mathfrak{F}$. Also, $\sigma({y})$ represents the $\sigma$-algebra generated by the random variable y{y}. For a function ${\varphi}\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$, define $\operatorname*{dom}{\varphi}\coloneqq\left\{{\mathsf{x}}\in\mathbb{R}^{d}:{\varphi}({\mathsf{x}})<+\infty\right\}$. $\varphi$} is proper if $\operatorname*{dom}\varphi\neq\varnothing$. The set of minimizers of $\varphi$} is $\operatorname*{arg\,min}{\varphi}=\{{\mathsf{x}}\in\mathbb{R}^{d}:{\varphi}({\mathsf{x}})=\inf{\varphi}\}$. If $\inf{\varphi}$ is finite, it is represented by $\varphi$$_{*}$. We denote the subdifferential of $\varphi$ at $\mathsf{x}$} as $\partial\varphi({\mathsf{x}})\coloneqq\left\{{{\mathsf{g}}}\in\mathbb{R}^{d}:{\varphi}({\mathsf{y}})-{\varphi}({\mathsf{x}})\geq\langle{{\mathsf{g}}},{\mathsf{y}}-{\mathsf{x}}\rangle\;(\forall\,{\mathsf{y}}\in\mathbb{R}^{d})\right\}$. When $\varphi$} is differentiable $\nabla{\varphi}$ denotes the gradient of $\varphi$}. $\varphi$} is said $L$-smooth if its gradient is $L$-Lipschitz continuous. The (convex) conjugate of $\varphi$} is the function ${\varphi}^{*}\colon\mathbb{R}^{d}\to[-\infty,+\infty]$ defined by ${\mathsf{y}}\mapsto\sup_{{\mathsf{x}}\in\mathbb{R}^{d}}\langle{\mathsf{y}},{\mathsf{x}}\rangle-{\varphi}({\mathsf{x}}).$ In this work, $\ell^{1}$ represents the space of sequences with summable norms and $\ell^{2}$ with summable squared norms. We set $\mathbb{R}^{d}_{+}=\{{\mathsf{x}}\in\mathbb{R}^{d}:\mathsf{x}_{i}\geq 0,i\in[d]\}$ and $\mathbb{R}^{d}_{++}=\{{\mathsf{x}}\in\mathbb{R}^{d}:\mathsf{x}_{i}>0,i\in[d]\}$.
@@ 2.2 Bregman distance
The Bregman distance $D_{h}({\mathsf{x}},{\mathsf{y}})$, between ${\mathsf{x}}\in\mathcal{C},{\mathsf{y}}\in\operatorname{int}\mathcal{C}$, is defined as $D_{h}({\mathsf{x}},{\mathsf{y}})\coloneqq h({\mathsf{x}})-h({\mathsf{y}})-\langle\nabla h({\mathsf{y}}),{\mathsf{x}}-{\mathsf{y}}\rangle$, where $h\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$, often called kernel, is a strictly convex and twice continuously differentiable function on $\operatorname{int}\mathcal{C}$, with $\operatorname*{dom}h=\mathcal{C}$ and $\lim_{{\mathsf{x}}\in\operatorname*{int}\mathcal{C},{\mathsf{x}}\to\text{bd}\,C}\|\nabla h(x)\|=+\infty$. In what follows, $h$ will always refer to a function with those properties, unless stated otherwise. A classical example is the usual squared Euclidean distance $D_{h}({\mathsf{x}},{\mathsf{y}})=\frac{1}{2}\|{\mathsf{x}}-{\mathsf{y}}\|_{2}^{2}$ when $h=\frac{1}{2}\|\cdot\|_{2}^{2}$. The Bregman distance often captures the underlying geometry of an optimization problem more effectively than the Euclidean distance. It can be better suited to the constraints set or to the objective function properties. This makes it a powerful tool in optimization. For instance, the Kullback–Leibler divergence on $\mathbb{R}^{d}$ does not have a global Lipschitz continuous gradient, however it is relatively smooth w.r.t. the kernel $h({\mathsf{x}})=-\sum_{i}\log\mathsf{x}_{i}$; see Bauschke et al. [2]. Unfortunately, the Bregman distance is not necessarily symmetric, homogeneous nor translation invariant. The latter two shortcomings are the reasons why the following assumption is required throughout the paper.
Assumption 2.1.
@@
$D_{h^{*}}({\mathsf{x}}+\lambda{\mathsf{z}},{\mathsf{x}})\leq G({\mathsf{x}},{\mathsf{y}},{\mathsf{z}})\lambda^{2}D_{h^{*}}({\mathsf{y}}+{\mathsf{z}},{\mathsf{y}}).$ (2.1)
According to Dragomir et al. [11], this assumption seems unavoidable when using past iterates in an algorithm with Bregman distance. This is also the case of accelerated methods where similar assumption is made; see Hanzely et al. [14]. The two papers provide extensive discussion and numerous examples regarding this assumption. Below, we present just one possible instance.
@@
    From Bauschke et al. [3, Theorem 3.7], we know that, for all $({\mathsf{x}},{\mathsf{y}})\in(\operatorname{int}\,\mathcal{C})^{2}$, $D_{h}({\mathsf{x}},{\mathsf{y}})=D_{h^{*}}(\nabla h({\mathsf{y}}),\nabla h({\mathsf{x}}))$ . Hence, $\gamma_{h}=\gamma_{h^{*}}$.
@@ 3 Problem setting
First, we (re)introduce the finite-sum problem and collect all assumptions that we need for this paper, while not all of them will be required to hold at the same time. Section 4() presents the proposed algorithm for solving such problems:
@@
    ${\mathsf{x}}_{*}\in\operatorname*{int}\mathcal{C}$ such that $\nabla F({\mathsf{x}}_{*})=0$ and $h^{*}$ is full domain.
Assumption (A.i) is always supposed true in this paper. Assumptions (A.ii) and (A.iv) extend the Euclidean notions of strong convexity and $L$-smoothness (i.e., $L$-Lipschitz continuity of the gradient), respectively, to the Bregman setting. Assumption (A.ii), with $\beta>0$, is only required when establishing a linear convergence rate (to a neighborhood of the optimum) for the vanilla BSPPA algorithm. Similarly, (A.iv), which corresponds to the relative strong convexity of $F$, will be used to derive linear convergence rates to the exact minimum for the variance-reduced variants of BSPPA; in that case, only convexity of each $f_{i}$ will be assumed. Finally, (A.iv) and (A.ii) with $\beta>0$ are not required to obtain the sublinear convergence rates. Regarding Assumption (A.v), exactly like [11], our analysis can not deal with the case where ${\mathsf{x}}_{*}\in\text{bd}\,\mathcal{C}$. It would be a valuable extension to our work.
@@ 4 BSPPA with generic variance reduction
We propose to solve Problem (3.1) using the following Bregman stochastic proximal point algorithm (BSPPA) with a generic variance reduction satisfying the abstract conditions in Assumptions 4.5() further below.
@@
We assume in all this work that this minimization has a unique solution in $\operatorname*{int}\mathcal{C}$. Compared to BSPPA, we have an additional linear perturbation that contains ${e}_{k}$. It is the generic variance reduction term. As we shall see in Section 5(), depending on the specific algorithms, ${e}_{k}$ may be defined in various ways. When ${e}_{k}$ is set to 0, BSPPA is recovered. Using the optimality condition, the update at iteration $k$ can be rewritten as:
@@
where ${g}_{k+1}\in\partial f_{i_{k}}({x}_{k+1})$ is such that (4.1) holds. Since ${w}_{k}$ depends on ${g}_{k+1}$, Equation (4.1) shows that Algorithm 4.1() is an implicit algorithm. This is in contrast to an explicit update
@@
and is assumed well-posed.
@@
The virtual explicit sequences $({v}_{k})_{k\in\mathbb{N}}$ and $({z}_{k})_{k\in\mathbb{N}}$ introduced in (4.2) and (4.3) are crucial to our analysis. They will be used to prove the main proposition even though they do not appear in Algorithm 4.1().
@@
The integration of these explicit iterates is the reason why our analysis extends seamlessly to the Bregman SGD case, providing a unified variance reduction study for Bregman SGD. All the variance reduction results will stand true for SGD with Bregman distance; see Remark B.2() and Proposition B.3().
The next assumptions that we consider concern the noise and variance of Algorithm 4.1(). The virtual explicit sequence $({z}_{k})_{k\in\mathbb{N}}$ appears in those assumptions for the sake of analysis. This allows, at the same time, the application of the analysis and the results to the Bregman SGD case.
@@
These assumptions on the noise and variance of the generic Algorithm 4.1() are “assumed” for the general analysis. But when algorithms are specified in Section 5() by defining ${e}_{k}$, they will be proved by corresponding lemmas. They read as follows.
@@
    $\mathsf{E}[{e}_{k}\,|\,\mathfrak{F}_{k}]=0$ and $\mathsf{E}[{g}_{k}\,|\,\mathfrak{F}_{k}]\in\partial F({x}_{k})$,
@@
    $\mathsf{E}[D_{h}({x}_{k},{z}_{k+1})\,|\,\mathfrak{F}_{k}]\leq\alpha_{k}^{2}A_{k}\left(F({x}_{k})-F({\mathsf{x}}_{*})\right)+\alpha_{k}^{2}B_{k}\sigma_{k}^{2}+\alpha_{k}^{2}N_{k}$,
@@
In the smooth case, Assumption (B.i) ensures that $\mathsf{E}[{v}_{k}\,|\,\mathfrak{F}_{k}]=\mathsf{E}[\nabla f_{i_{k}}({x}_{k})\,|\,\mathfrak{F}_{k}]=\nabla F({x}_{k})$, so that the direction ${v}_{k}$ is an unbiased estimator of the full gradient of $F$ at ${x}_{k}$, which is a standard assumption in the related literature. Assumption (B.ii) on $\mathsf{E}\left[D_{h}({x}_{k},{z}_{k+1})\,|\,\mathfrak{F}_{k}\right]$ is the equivalent of what is called, in the literature [17, 10] and in the Euclidean setting, the expected smoothness or $ABC$-assumption on $\mathsf{E}\left[\lVert\nabla f_{i_{k}}({x}_{k})\rVert^{2}\,|\,\mathfrak{F}_{k}\right]$ with $\sigma_{k}=\|\nabla F({x}_{k})\|$ and $N_{k}$ constant. Assumption (B.iii) provides some control on the variance from iteration to iteration. Indeed, as we will see later that the sequence $(\sigma_{k})_{k\in\mathbb{N}}$ encodes the variance of the Algorithm 4.1(). Depending on $\rho$, (B.iii) makes sure that the variance does not blow up and possibly reduces along the iterations whenever the algorithm converges.